OpenStudy (anonymous):

What is the oblique asymptote of the function f(x)= [x^2+5x+6]/[x+4]?

OpenStudy (anonymous):

To find the slant asymptote, do the long division...The slant asymptote is the polynomial part of the answer, not the remainder.

OpenStudy (anonymous):

(1+x)?

OpenStudy (anonymous):

| | x | + | 4 | | x | + | 1 x^2 | + | 5 x | + | 6 x^2 | + | 4 x | | | | x | + | 6 | | x | + | 4 | | | | 2

OpenStudy (anonymous):

thats right :)

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

so we set x+1 = 0

OpenStudy (anonymous):

oblique asymptote is a line so y=x+1

OpenStudy (anonymous):

[x^2-x-12]/[x-4] are these correct: disc.= -3 x int=(3,0) y int= (0,-3) shape= straight line